# Semi-Blind Strategies for MMSE Channel Estimation Utilizing Generative Priors

### v1

@@ Abstract
This paper investigates semi-blind channel estimation for massive multiple-input multiple-output (MIMO) systems. To this end, we first estimate a subspace based on all received symbols (pilot and payload) to provide additional information for subsequent channel estimation. We show how this additional information enhances minimum mean square error (MMSE) channel estimation. Two variants of the linear MMSE (LMMSE) estimator are formulated, where the first one solves the estimation within the subspace, and the second one uses a subspace projection as a preprocessing step. Theoretical derivations show the superior estimation performance of the latter method in terms of mean square error for uncorrelated Rayleigh fading. Subsequently, we introduce parameterizations of this semi-blind LMMSE estimator based on two different conditional Gaussian latent models, i.e., the Gaussian mixture model and the variational autoencoder. Both models learn the underlying channel distribution of the propagation environment based on training data and serve as generative priors for semi-blind channel estimation. Extensive simulations for real-world measurement data and spatial channel models show the superior performance of the proposed methods compared to state-of-the-art semi-blind channel estimators with respect to the MSE.
Index Terms:
Semi-blind channel estimation, Gaussian mixture model, variational autoencoder, measurement data.
@@ I Introduction
Accurate channel state information (CSI) is crucial for achieving the expected high data rates promised by multiple-input-multiple-output (MIMO) systems [2, 3, 4]. The CSI describes the communication link between transmitter and receiver, characterized by its time-varying and frequency-selective nature, which is prone to rapid changes making the task of channel estimation complex [5]. As accurate channel estimates are essential for the successful transmission of data, it is placed at the center of several research efforts [6, 7].
@@
The most widely adopted methods in wireless communication utilize known training or pilot symbols transmitted across the channel using some of the radio resource blocks [8]. Afterward, the receiver uses the observed signals to determine a reliable CSI estimate. As the number of pilots scales with the number of users, the spectral efficiency decreases for higher number of users as less symbols are available for transmitting data. To enhance channel estimation without increasing the number of pilot symbols, various methods have been developed that leverage the information embedded in the observed data symbols at the receiver to infer channel characteristics [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. These methods exploit structure and redundancy within the transmitted data and yield more accurate CSI estimates.
@@
The benefit of semi-blind channel estimation was first studied in [9], where Cramer-Rao bounds (CRBs) for blind, semi-blind, and training-based channel estimation were investigated in the context of single-input-multiple-output (SIMO) systems. In [10, 11], semi-blind channel estimation schemes based on maximum likelihood (ML) estimation were introduced. The asymptotic performance of the respective estimators was also studied in [10] for infinitely long data sequences. Another asymptotic behavior, where the number of antennas grows to infinity, was studied in [12]. Here, the authors identified two interference components in semi-blind channel estimation, which do not vanish even for large numbers of antennas. Early work on improving the least squares (LS) estimator using a semi-blind algorithm was conducted in [13]. The authors in [14] propose to use partially decoded data for channel estimation. Furthermore, in [15], semi-blind and blind channel estimation was studied to enhance the maximum a-posteriori (MAP) channel estimates in massive MIMO systems. These findings are based on favorable propagation, which only holds for large antenna arrays deployed at the base station (BS). In [16], two semi-blind channel estimators based on the expectation maximization (EM) algorithm are studied. The assumption of a Gaussian distribution for the data was verified, leading to a closed-form solution for the E-step. Another iterative framework optimizing the likelihood based on message passing (MP) is used in [17, 18] and references therein. In [19], a data-aided iterative scheme is proposed for orthogonal time frequency space (OTFS) systems by employing affine-precoded superimposed pilots. The performance improvement achieved with these iterative approaches generally requires computational costly updates. A low complexity iterative LS channel estimation algorithm is proposed in [20] for a massive MIMO turbo-receiver. In [21], the concept of semi-blind channel estimation was adapted for time domain synchronous-orthogonal frequency division multiplexing (OFDM) systems, where in addition to the pseudo noise sequence in the guard interval, the sent OFDM data symbols are exploited for the channel estimation. In [22], a framework was introduced for choosing reliable decoded data symbols, which can be interpreted as additional pilots. Similarly, reliably detected symbols are used for a semi-data-aided channel estimation in [23]. In [24], peak-power carriers in an OFDM system are selected to eliminate the need to determine reliable data symbols at the receiver. Recently, a diffusion model-based approach for joint channel estimation and detection was proposed in [25], where a diffusion process is constructed that models the joint distribution of the channels and symbols given noisy observations.
@@
In this article, we focus on pilot-based estimators which minimize the MSE and investigate how these estimators can be extended for the semi-blind case. Notably, in [26] a subspace formulation of the MMSE estimator was used to mitigate pilot contamination in massive MIMO systems. The MMSE estimator is known to be the conditional mean estimator (CME) [27, Ch. 10], which, in general, is intractable and can not be computed in closed form. Recently, powerful approximations based on machine learning were presented in [28, 29, 30, 31, 32]. The benefit of machine learning is to enhance the task at hand by using prior information captured during the learning stage. For a given BS cell environment, the probability density function (PDF) representing potential user channels can be considered valuable prior information. Since this true underlying distribution is unknown, machine learning methods rely on a representative data set, which is assumed to be available at the BS. Based on this data set the PDF of the user channels can be learned. The first proposal of using a Gaussian mixture model (GMM) to formulate an estimator was done in [28] for the case of image processing. The approaches in [30, 31, 32] build on that by constructing a conditionally Gaussian latent model (CGLM) for the PDF of a BS cell environment. The learned CGLM not only enables MMSE channel estimation in [30, 31, 32] but can also be used for e.g., a limited feedback scheme as in [33]. In this work, we propose to utilize CGLMs to parameterize the CME in the semi-blind setting.
The contributions of this work are summarized as follows:
@@
  • We introduce two variants of the linear MMSE (LMMSE) estimator incorporating subspace knowledge provided by the payload data symbols. First, we depict how the LMMSE channel estimator can solve a subspace estimation problem [26]. As an alternative, we propose a projection method that is computationally more efficient since it allows for the pre-calculation of LMMSE filters.
@@
  • With theoretical derivations we show the superior MSE performance of the proposed projection method in the case of uncorrelated Rayleigh fading and perfect subspace knowledge.
@@
  • Extensive simulations on different datasets, consisting of typical massive MIMO systems with multiple users and including real-world measurement data, show the superior performance of our proposed methods compared to state-of-the-art semi-blind channel estimation algorithms with respect to the MSE.
Preliminary results were presented in [1] and extended to the multi-user MIMO case in [34], which we extend further in the following aspects. The theoretical analyses in Section III enhance the foundation of the proposed semi-blind channel estimation strategies and provide analytic insights into the superior performance of the proposed projection method. We extend our concept of semi-blind MMSE channel estimation to the whole class of CGLMs, providing a more general framework to parameterize the semi-blind LMMSE estimator. Finally, we provide more comprehensive simulation results to show the strengths of our proposed strategies.
[…]
@@ II System and Channel Model
We consider a multi-user uplink system with $J$ single-antenna users and a BS equipped with $M$ receive antennas. The received signal vector at time instance $n$ is then
@@
where ${\bm{x}}(n)=[x_{1}(n),...,x_{J}(n)]^{\mathrm{T}}\in\mathbb{C}^{J}$ and ${\bm{n}}(n)\in\mathbb{C}^{M}$ denote the signal sent by each of the $J$ users and the additive noise, respectively, whereas ${\bm{H}}=[{\bm{h}}_{1},...,{\bm{h}}_{J}]$ contains the individual channels of the users ${\bm{h}}_{j}\in\mathbb{C}^{M}$. The case of multiple antennas at the users can be transformed into (1) by viewing each active stream as a different user. The corresponding channel ${\bm{h}}_{j}$ can then be seen as the effective channel. For further details on semi-blind channel estimation in multi-user MIMO, we refer the reader to [34]. For the task of channel estimation, we consider a channel coherence interval larger than the number of snapshots $N$, i.e., the channels are constant over all snapshots. We assume that the noise is Gaussian with ${\bm{n}}(n)\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M})$.
@@
where ${\bm{Y}}\in\mathbb{C}^{M\times N}$, ${\bm{Y}}^{\prime}_{p}\in\mathbb{C}^{M\times N_{p}}$, ${\bm{Y}}_{d}\in\mathbb{C}^{M\times N-N_{p}}$, ${\bm{P}}\in\mathbb{C}^{J\times N_{p}}$, and ${\bm{D}}\in\mathbb{C}^{J\times N-N_{p}}$ denote all received observations, received pilot observations, received payload data observations, sent pilots, and sent payload data symbols, respectively. In order to fully illuminate the channels, the number of pilots is, at minimum, the number of users $N_{p}\geq J$, and orthogonal pilots are used. We set $N_{p}=J$, and utilize discrete Fourier transform (DFT) pilot sequences. After decorrelating the orthogonal pilot sequences the received pilot observations simplify to
${\bm{Y}}_{p}={\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}={\bm{H}}{\bm{P}}{\bm{P}}^{\mathrm{H}}+{\bm{N}}{\bm{P}}^{\mathrm{H}}={\bm{H}}+{\bm{N}}_{p},$ (3)
where ${\bm{N}}_{p}$ has the same statistics as ${\bm{N}}$ and, hence, we can omit the subscript. This enables to consider channel estimation from a per user perspective in the subsequent discussions. For reasons of simpler readability, the index for the respective user is, therefore, no longer given in the following. Consequently, we denote the pilot observation of a user as
@@
with ${\bm{n}}\sim\mathcal{N}_{\mathbb{C}}(0,{\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M})$.
@@ II-A Spatial Channel Model
based on a set of parameters ${\bm{\delta}}$, which describe the directions and properties of the multi-path propagation clusters. The main angles are drawn independently from a uniform distribution in $[0,2\pi]$ and the path gains are independent zero-mean Gaussians. The spatial covariance matrix is given by
@@
where $g(\vartheta,{\bm{\delta}})$ is the power density consisting of a weighted sum of Laplace densities, which have standard deviations $\sigma_{\text{AS}}$ corresponding to the angular spread of the propagation clusters. The BS employs a uniform linear array (ULA) with $M=64$ antennas and $\lambda/2$ spacing. The steering vector is then given as
${\bm{a}}(\vartheta)=\frac{1}{\sqrt{M}}\left[1,\mathrm{e}^{-\mathrm{j}\pi\sin(\vartheta)},\dots,\mathrm{e}^{-\mathrm{j}\pi(M-1)\sin(\vartheta)}\right]^{\mathrm{T}}.$ (7)
For every channel sample, we consider a new ${\bm{\delta}}$ and draw the sample according to ${\bm{h}}\sim\mathcal{N}_{\mathbb{C}}\left(\bm{0},{\bm{C}}_{\bm{\delta}}\right)$.

### v2

@@ Abstract
This paper investigates semi-blind channel estimation for massive multiple-input multiple-output (MIMO) systems. To this end, we first estimate a subspace based on all received symbols (pilot and payload) to provide additional information for subsequent channel estimation. This additional information enhances minimum mean square error (MMSE) channel estimation. Two variants of the linear MMSE (LMMSE) estimator are formulated, where the first one solves the estimation within the subspace, and the second one uses a subspace projection as a preprocessing step. Theoretical derivations show the latter method’s superior estimation performance in terms of mean square error for uncorrelated Rayleigh fading. Further, we provide asymptotical insights on how the proposed MMSE-based channel estimation strategy outperforms the unbiased Cramer-Rao bound. Subsequently, we introduce parameterizations of these semi-blind LMMSE estimators based on two different conditional Gaussian latent models, i.e., the Gaussian mixture model and the variational autoencoder. Both models learn the propagation environment’s underlying channel distribution based on training data and serve as generative priors for our semi-blind channel estimation. Extensive simulations for real-world measurement data and spatial channel models show the proposed methods’ superior performance compared to state-of-the-art semi-blind channel estimators in terms of MSE.
@@ I Introduction
Accurate channel state information (CSI) is crucial for achieving the expected high data rates promised by multiple-input-multiple-output (MIMO) systems [2, 3, 4]. The CSI describes the communication link between transmitter and receiver, characterized by its time-varying and frequency-selective nature, which is prone to rapid changes, making the task of channel estimation complex [5]. As accurate channel estimates are essential for successfully transmitting data, it is at the center of several research efforts [6, 7].
@@
The most widely adopted methods in wireless communications utilize known training or pilot symbols transmitted across the channel using some of the radio resource blocks [8]. Afterward, the receiver uses the observed signals to determine a reliable CSI estimate. As the number of pilots scales with the number of users, the spectral efficiency decreases for a higher number of users as fewer symbols are available for transmitting data. To enhance channel estimation without increasing the number of pilot symbols, various methods have been developed that leverage the information embedded in the observed data symbols at the receiver to infer channel characteristics [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. These methods exploit structure and redundancy within the transmitted data and yield more accurate CSI estimates.
@@
The benefit of semi-blind channel estimation was first studied in [9], where Cramer-Rao bounds (CRBs) for blind, semi-blind, and training-based channel estimation were investigated in the context of single-input-multiple-output (SIMO) systems. In [10, 11], semi-blind channel estimation schemes were introduced based on maximum likelihood (ML) estimation. The respective estimators’ asymptotic performances were also studied in [10] for infinitely long data sequences. Another asymptotic behavior, where the number of antennas grows to infinity, was studied in [12]. Here, the authors identified two interference components in semi-blind channel estimation, which do not vanish even for large numbers of antennas. Early work on improving the least squares (LS) estimator using a semi-blind algorithm was conducted in [13]. Furthermore, in [14], semi-blind and blind channel estimation was studied to enhance the maximum a-posteriori (MAP) channel estimates in massive MIMO systems. These findings are based on favorable propagation, which only holds for large antenna arrays deployed at the base station (BS). In [15], two semi-blind channel estimators are studied based on the expectation maximization (EM) algorithm. The assumption of a Gaussian distribution for the data was verified, leading to a closed-form solution for the E-step. Another iterative framework optimizing the likelihood based on message passing (MP) is used in [16, 17] and references therein. In [18], a data-aided iterative scheme is proposed for orthogonal time frequency space (OTFS) systems by employing affine-precoded superimposed pilots. The performance improvement achieved with these iterative approaches generally requires computationally costly updates. A low complexity iterative LS channel estimation algorithm is proposed in [19] for a massive MIMO turbo-receiver. In [20], the concept of semi-blind channel estimation was adapted for time domain synchronous-orthogonal frequency division multiplexing (OFDM) systems, where in addition to the pseudo noise sequence in the guard interval, the sent OFDM data symbols are exploited for the channel estimation. In [21, 22, 23], different decision-directed frameworks were proposed, which treat reliably decoded data symbols as additional pilots. All of these methods utilize iterative approaches, thus exhibiting large complexities. In [24], peak-power carriers in an OFDM system are selected to eliminate the need to determine reliable data symbols at the receiver. Recently, a diffusion model-based approach for joint channel estimation and detection was proposed in [25], where a diffusion process is constructed that models the joint distribution of the channels and symbols given noisy observations.
@@
Many state-of-the-art estimators fall in the class of unbiased estimators, which are potentially MSE-suboptimal. Instead, in this article, we focus on general, potentially biased, pilot-based estimators that minimize the MSE by utilizing prior information about the channel distribution, and investigate how these estimators can be extended for the semi-blind case. Contrarily to previous works, e.g., [21, 22, 23], we do not focus on using decoded data symbols as additional pilot but instead want to improve the estimators themselves. Although both approaches can be combined. Notably, in [26] a subspace formulation of the MMSE estimator was already used to mitigate pilot contamination in massive MIMO systems. The MMSE estimator is known to be the conditional mean estimator (CME) [27, Ch. 10], which, in general, is intractable and can not be computed in closed form. Recently, powerful approximations based on machine learning were presented in [28, 29, 30, 31, 32]. The benefit of machine learning is to enhance the task at hand by using prior information captured during the learning stage. For a given BS cell environment, the probability density function (PDF) representing potential user channels can be considered valuable prior information. Since this true underlying distribution is unknown, machine learning methods rely on a representative data set, which is assumed to be available at the BS. Based on this data set, the user channels’ PDF can be learned. The first proposal of using a Gaussian mixture model (GMM) to formulate an estimator was done in [28] for the case of image processing. The approaches in [30, 31, 32] build on that by constructing a conditionally Gaussian latent model (CGLM) for the PDF of a BS cell environment. The learned CGLM not only enables MMSE channel estimation in [30, 31, 32] but can also be used for, e.g., a limited feedback scheme as in [33]. In this work, we propose to utilize CGLMs to parameterize the CME in the semi-blind setting.
This work’s contributions are summarized as follows:
@@
  • We introduce two variants of the linear MMSE (MMSE) estimator incorporating subspace knowledge provided by the payload data symbols. Firstly, we depict how the LMMSE channel estimator can solve a subspace estimation problem [26]. As an alternative, we propose a new projection method that is computationally more efficient since it allows for the pre-calculation of LMMSE filters.
@@
  • With theoretical derivations, we show the proposed projection method’s superior MSE performance in the case of uncorrelated Rayleigh fading and perfect subspace knowledge.
  • We provide an asymptotical analysis of how the projected LMMSE outperforms the unbiased CRB for perfect subspaces. These derivations highlight the MMSE-based channel estimation’s superiority compared to unbiased channel estimation in the case of semi-blind channel estimation.
@@
  • Extensive simulations on different datasets, consisting of typical massive MIMO systems with multiple users and real-world measurement data, show our proposed methods’ superior performance compared to state-of-the-art semi-blind channel estimation algorithms in terms of the MSE.
Preliminary results were presented in [1] and extended to the multi-user MIMO case in [34], which we extend further in the following aspects. The theoretical analyses in Section III enhance the proposed semi-blind channel estimation strategies’ foundation and provide analytic insights into the proposed projection method’s superior performance. We extend our concept of semi-blind MMSE channel estimation to the whole class of CGLMs, providing a more general framework to parameterize the semi-blind LMMSE estimator. Finally, we provide more comprehensive simulation results to show our proposed strategies’ strengths.
Notations: Matrices and vectors are denoted with boldface symbols. $\bm{0}$ and $\mathbf{I}_{N}$ denote the zero vector of appropriate size and the identity matrix of size $N\times N$, respectively. $\mathbb{E}[\cdot]$, $\mathrm{tr}(\cdot)$, and $\mathrm{range}(\cdot)$ denote the expectation, trace, and range operators, respectively. We use $(\cdot)^{\mathrm{T}}$, $(\cdot)^{\mathrm{H}}$, $(\cdot)^{-1}$, $(\cdot)^{\dagger}$ to denote the transpose, conjugate transpose, inverse, and pseudo inverse. $\|\cdot\|$ denotes the $2$-norm of a vector. $\mathcal{N}_{\mathbb{C}}(\bm{\mu},{\bm{C}})$ denotes the circularly symmetric complex Gaussian distribution with mean $\bm{\mu}$ and covariance matrix $\bm{C}$}.
@@ II System and Channel Model
We consider a multi-user system with $J$ single-antenna users and a BS equipped with $M$ receive antennas. We focus on the uplink scenario for this work, but the proposed methods can also be extended to the downlink. The received signal vector at time instance $n$ is then
@@
where ${\bm{x}}(n)=[x_{1}(n),...,x_{J}(n)]^{\mathrm{T}}\in\mathbb{C}^{J}$ and ${\bm{n}}(n)\in\mathbb{C}^{M}$ denote the signal sent by each of the $J$ users and the additive noise, respectively, whereas ${\bm{H}}=[{\bm{h}}_{1},...,{\bm{h}}_{J}]$ contains the individual user channels ${\bm{h}}_{j}\in\mathbb{C}^{M}$. The case of multiple antennas at the users can be transformed into (1) by viewing each active stream as a different user with its corresponding channel ${\bm{h}}_{j}$ being the respective effective channel. For further details on semi-blind channel estimation in multi-user MIMO, we refer the reader to [34]. For the case of downlink semi-blind channel estimation, a subspace of the effective channel needs to be identifiable to improve estimation performance. This is the case if not all degrees of freedom are utilized, e.g., fewer pilots than receive antennas. We consider a channel coherence interval larger than the number of snapshots $N$, i.e., the channels are constant over all snapshots. We assume that the noise is Gaussian with ${\bm{n}}(n)\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M})$ and the sent signals satisfy $\mathbb{E}[{\bm{x}}(n){\bm{x}}^{\mathrm{H}}(n)]=\frac{1}{J}\mathbf{I}_{J}$.
@@
where ${\bm{Y}}\in\mathbb{C}^{M\times N}$, ${\bm{Y}}^{\prime}_{p}\in\mathbb{C}^{M\times N_{p}}$, ${\bm{Y}}_{d}\in\mathbb{C}^{M\times N-N_{p}}$, ${\bm{P}}\in\mathbb{C}^{J\times N_{p}}$, and ${\bm{D}}=[{\bm{d}}_{1},\dots,{\bm{d}}_{J}]^{\mathrm{T}}$ denote all received observations, received pilot observations, received payload data observations, sent pilots, and sent payload data symbols, respectively. The $j$-th user’s payload symbols ${\bm{d}}_{j}\in\mathbb{S}_{j}^{N-N_{p}}$ include realizations of the used transmit constellation, where $\mathbb{S}_{j}$ denotes the set of constellation points of user $j$. In general, the employed user constellations can differ from each other. In order to fully illuminate the channels, the number of pilots is, at minimum, the number of users $N_{p}\geq J$, and orthogonal pilots are used. We set ${\bm{P}}{\bm{P}}^{\mathrm{H}}=\frac{N_{p}}{J}\mathbf{I}_{M}$, and utilize submatrices of the discrete Fourier transform (DFT) matrix of size $N_{p}\times N_{p}$. After decorrelating the orthogonal pilot sequences, the received pilot observations simplify to
${\bm{Y}}_{p}={\bm{Y}}^{\prime}_{p}{\bm{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}={\bm{H}}{\bm{P}}{\bm{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}+{\bm{N}}{\bm{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}={\bm{H}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\frac{J}{N_{p}}}}{\bm{N}}_{p},$ (3)
where ${\bm{N}}_{p}$ has the same statistics as $\bm{N}$} and, hence, we can omit the subscript. This decorrelation lets us consider channel estimation from a per-user perspective in the subsequent discussions. For reasons of simpler readability, the index for the respective user is, therefore, no longer given in the following. Consequently, we denote a user’s pilot observation as
@@
with ${\bm{n}}\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{\bm{n}}=\sigma^{2}_{\mathrm{eff}}\mathbf{I}_{M})$, where $\sigma^{2}_{\mathrm{eff}}=\sigma^{2}\frac{J}{N_{p}}$ denotes the effective noise variance. Further, we assume normalized channels as $\mathbb{E}[\|{\bm{h}}\|^{2}]=M$, which lets us define the signal-to-noise ratio (SNR) as $\sigma^{-2}$. As the number of pilots $N_{p}$ results in a simple scaling of SNR in the pilot transmission phase, we use $N_{p}=J$ in the rest of this work.
@@ II-A Spatial Channel Model
based on a set of parameters $\bm{\delta}$}, which describe the directions and properties of the multi-path propagation clusters. The central angles are drawn independently from a uniform distribution in $0,2\pi]$, and the path gains are independent zero-mean Gaussians. The spatial covariance matrix is given by
@@
where $g(\vartheta,{\bm{\delta}})$ is the power density consisting of a weighted sum of Laplace densities, which have standard deviations $\sigma_{\text{AS}}$ corresponding to the propagation clusters’ angular spread. The BS employs a uniform linear array (ULA) with $M$ antennas and $\lambda/2$ spacing. The steering vector is then given as
${\bm{a}}(\vartheta)=\frac{1}{\sqrt{M}}\left[1,\mathrm{e}^{-\mathrm{j}\pi\sin(\vartheta)},...,\mathrm{e}^{-\mathrm{j}\pi(M-1)\sin(\vartheta)}\right]^{\mathrm{T}}.$ (7)
We consider a new $\bm{\delta}$} for every channel sample and draw the sample according to ${\bm{h}}\sim\mathcal{N}_{\mathbb{C}}\left(\bm{0},{\bm{C}}_{\bm{\delta}}\right)$.